OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x-8 A) y=1/2x+7 B) y=1/2x+5 C) y=2+7 D) y=2x+5

OpenStudy (precal):

parallel to the equation, what doe that mean in terms of the slope?

OpenStudy (precal):

are you looking for the same slope or opposite reciprocals? If you know the answer to that question, then you can eliminate 2 of your choices.

OpenStudy (anonymous):

what?....

OpenStudy (precal):

 your equation y=(1/2)x-8 looks like this

OpenStudy (precal):

 parallel lines look like this, in your equation one of the numbers guarantee that

OpenStudy (precal):

Ok I will let rox13h help you............ Process is more important than solutions, if you know the process then you will be able to solve all problems of this form...

OpenStudy (anonymous):

if it's parallel, then the slope should be the same, which is 1/2 that leaves a and b... now put the coordinates (-2,6) to both equations and the equation that satisfies it is the answer. the answer is a because y=1/2x +7 6+1=7